# Decoupling and randomization for double-indexed permutation statistics

### v1

@@ 1 Introduction
It is known that the double-indexeded sum in (1.1) admits a decomposition into a simple combinatorial sum and another double-indexed sum induced by a degenerate tensor $\mathbf{D}=[d(i,j,k,\ell)]\in\mathbb{R}^{N\times N\times N\times N}$. Specifically, define
@@
In this paper, we develop new decoupling and randomization inequalities to simplify the task of bounding the moment-generating function (MGF) $\mathrm{E}\big[\exp(\lambda Q_{d})\big]$. Based on the bounds on the MGF of $Q_{d}$, we derive concentration inequalities for $Q_{d}$, and consequently also for the original statistic $Q_{w}$.
@@
where $\tau$ is an independent copy of $\pi$. The decoupling inequalities of interest compare $\mathrm{E}[\exp(\lambda Q_{d})]$ with (a variant of) $\mathrm{E}[\exp(\lambda\widetilde{Q}_{d})]$, the latter being typically much easier to analyze.
@@ 1.1 Motivation
Statistics of the form (1.1) are prevalent in nonparametric statistics and have also received increasing attention in causal inference. In nonparametric statistics, it has been noted that many quantities of statistical interest—including the correlation coefficients of Pearson, Kendall, Spearman, and more recently Chatterjee (Chatterjee, 2021), as well as discrepancy measures such as the Mann-Whitney-Wilcoxon statistic—can be expressed in this form and thus benefit from a unified treatment. In survey sampling and causal inference, such statistics have also become increasingly important within the framework of finite-population inference (Li and Ding, 2017; Ding, 2024; Han, 2024), where they play a role analogous to that of U-statistics in the super-population framework.
In the probability and statistical theory literature, quantities of the form (1.1) are often referred to as DIPS (Zhao et al., 1997), and may be viewed as natural higher-order extensions of Hoeffding’s statistics in the context of sampling without replacement (Hoeffding, 1951). The latter corresponds to the special—and in fact single-indexed—case $d(i,j,k,\ell)=a(i,k)$ for some second-order tensor $\mathbf{A}=[a(i,k)]\in\mathbb{R}^{N\times N}$.
@@ 1.2 Generalized combinatorial quadratic forms
Our version of decoupling and randomization enables the following tail bound for the generalized combinatorial quadratic form:
@@ 1.3 Examples in nonparametric statistics
$\mathrm{P}\Big(Q_{\sf U}-\mathrm{E}[Q_{\sf U}]>t\Big)\leq\exp\Big(-\frac{K\,t^{2}}{\,Nmn+Nt\,}\Big)\;+\;\exp\Big(-\frac{K\,t^{2}}{\,m^{2}n^{2}/N^{2}+(mn)^{1/2}t\,}\Big).$
@@
it is clear that $R$ has the same distribution as
@@
This framework encompasses several classical statistics as special cases, including Pearson’s correlation, Spearman’s rank correlation $\rho$, and Kendall’s $\tau$, under suitable choices of score functions:
@@
Chatterjee’s rank correlation was recently introduced by Chatterjee (2021) as an appealing alternative to the correlation coefficients in Example 1.2; see, e.g., Dette et al. (2013), Han (2021), Chatterjee (2024), and Lin and Han (2022) for more discussions. Under (1.3), Chatterjee’s rank correlation has the same distribution as
@@
$\mathrm{P}\big(\xi_{N}-\mathrm{E}[\xi_{N}]\geq t\big)\leq\exp\Big(-\frac{KN^{2}t^{2}}{1+Nt}\Big)\;+\;\exp\Big(-\frac{KNt^{2}}{1+\sqrt{N}\,t}\Big).$
@@ 1.4 An example in causal inference
Regression adjustment is widely used to improve efficiency in randomized experiments. Lei and Ding (2021) studied this method in completely randomized experiments with a diverging number of covariates and derived an exact finite-population decomposition of the regression-adjusted estimator $\widehat{\tau}_{\mathrm{adj}}$ proposed by Lin (2013). In the decomposition, certain finite-population bias terms appear in quadratic forms.
@@
be the set of units assigned to the treatment level $\omega$. The finite-population bias of $\widehat{\tau}_{\mathrm{adj}}$ can then be written as a function of
@@
Following Lei and Ding (2021), we assume $\mathbf{X}$ to be column-centered, which ensures that both $\mathbf{H}$ and $\mathbf{Q}(\omega)$ are doubly centered. Applying Theorem 1.1, we can show that there exists a universal constant $K>0$ such that, for any $t>0$,
@@ 2 Main results
$B:=\max_{\sigma\in\mathcal{S}_{N}}\Big\|[d(i,j,\sigma(i),\sigma(j))]_{i,j}\Big\|_{\sf op}~~{\rm and}~~V:=\frac{1}{N}\sum_{i,k\in[N]}\xi(i,k)^{2}+\frac{1}{N^{2}}\sum_{i,j,k,\ell\in[N]}d(i,j,k,\ell)^{2},$ (2.5)
@@
Recall from (1.2) that
$Q_{d}=\sum_{i,j\in[N]}d(i,j,\pi(i),\pi(j)).$
Then, for all $\lambda\in[0,1/5400B]$,
@@
Let $N\geq 20$ and $\mathbf{W}=[w(i,j,k,\ell)]\in\mathbb{R}^{N\times N\times N\times N}$ be a fixed tensor. Recall from (1.1) that
$Q_{w}=\sum_{i,j\in[N]}w(i,j,\pi(i),\pi(j)).$
Then, there exists a universal constant $K>0$ such that for any $t>0$,
@@
$\xi_{w}(i,k)=d_{w}(i,i,k,k)-\frac{1}{N}\sum_{k^{\prime}=1}^{N}d_{w}(i,i,k^{\prime},k^{\prime})-\frac{1}{N}\sum_{i^{\prime}=1}^{N}d_{w}(i^{\prime},i^{\prime},k,k)+\frac{1}{N^{2}}\sum_{i^{\prime},k^{\prime}\in[N]}d_{w}(i^{\prime},i^{\prime},k^{\prime},k^{\prime}),$
${\rm and}~~V_{d}:=\frac{1}{N}\sum_{i,k\in[N]}\xi_{w}(i,k)^{2}+\frac{1}{N^{2}}\sum_{i,j,k,\ell\in[N]}d_{w}(i,j,k,\ell)^{2}.$
Here $a_{w}(\cdot,\cdot)$ and $d_{w}(\cdot,\cdot,\cdot,\cdot)$ are introduced in (2) and (2), respectively.
@@ 3 Proofs of the main results
In the following, recall that $G_{1},G_{2},\ldots,G_{1}^{\prime},G_{2}^{\prime},\ldots,$ are independent standard Gaussian random variables, further independent of the system. Recall that $\bigl[\cdot\bigr]_{i,j}$ denotes the matrix whose $(i,j)$-th entry is given by the enclosed expression, where $i$ and $j$ range over the prescribed index sets (not necessarily $[N]$). For a set $I\subseteq[N]$, we write $|I|$ for its cardinality and $I^{c}:=[N]\setminus I$ for its complement. Additionally, for a permutation (or, more generally, a mapping) $\pi$, we denote $\pi(I):=\{\pi(i)\mid i\in I\}$.
@@ 3.1 Auxiliary lemmas
Consider $\mathbf{A}=[a(i,j)]\in\mathbb{R}^{N\times N}$ to be a (possibly random) matrix independent of $\pi$ and satisfying $\sum_{i,j\in[N]}a(i,j)=0$. We then have, for any $\lambda\in\mathbb{R}$,
@@
Consider $\mathbf{A}=[a(i,j)]\in\mathbb{R}^{N\times N}$ to be a fixed matrix. We then have, for any $|\lambda|<\sqrt{2}/B_{A}$,
@@
Consider $\mathbf{A}=[a(i,j)]\in\mathbb{R}^{N\times N}$ to be a fixed matrix whose entries are in $[0,1]$. Denote
$f(\pi)=\sum_{i=1}^{N}a(i,\pi(i)).$ (3.3)
Then for any $\lambda\geq 0$,
@@
Consider $\mathbf{A}=[a(i,j)]\in\mathbb{R}^{N\times N}$ to be a fixed matrix whose entries are in [1,1][-1,1]. Then for any $|\lambda|\leq 1/3$,
@@
Fix $\mathbf{A}$ as in Lemma 3.6. Then for any $\lambda\geq 0$,
@@
$V_{2}=\max_{I,J}\frac{1}{N_{1}(N-N_{1})}\sum_{\begin{subarray}{c}i,k\in[N]\\ j,\ell\in[M]\end{subarray}}\widetilde{d}_{I,J}(i,j,k,\ell)^{2}\leq\max_{I,J}\frac{1}{N_{1}(N-N_{1})}\sum_{\begin{subarray}{c}i,k\in[N]\\ j,\ell\in[M]\end{subarray}}d_{I,J}(i,j,k,\ell)^{2}\leq 4.1V$
@@
$\xi_{w}(i,k)$ $=d_{w}(i,i,k,k)-\frac{1}{N}\sum_{k^{\prime}=1}^{N}d_{w}(i,i,k^{\prime},k^{\prime})-\frac{1}{N}\sum_{i^{\prime}=1}^{N}d_{w}(i^{\prime},i^{\prime},k,k)-\frac{1}{N^{2}}\sum_{i^{\prime},k^{\prime}\in[N]}d_{w}(i^{\prime},i^{\prime},k^{\prime},k^{\prime})$
$=\Big(c_{ii}-\frac{1}{N}\sum_{i^{\prime}=1}^{N}c_{i^{\prime}i^{\prime}}\Big)\Big(\widetilde{a}_{kk}+\frac{1}{N}\sum_{k^{\prime}=1}^{N}\widetilde{a}_{k^{\prime}k^{\prime}}\Big).$
@@ 4 Proofs of auxiliary lemmas
By Chatterjee’s third lemma (Han, 2024, Lemma 3.11), taking $\psi=\frac{3}{2}\lambda^{2}$, we have
@@
where $r(\cdot)$ is defined as in (Han, 2024, Lemma 3.11); for completeness, we recall that for any $\psi>0$,
@@
$f^{\downarrow}(\pi)=\sum_{i=1}^{N}a(i,\pi(i))\cdot\mathds{1}(|a(i,\pi(i))|\leq\rho)~~{\rm and}~~f^{\uparrow}(\pi)=\sum_{i=1}^{N}|a(i,\pi(i))|\cdot\mathds{1}(|a(i,\pi(i))|>\rho),$

### v2

@@ 1 Introduction
It is known that the double-indexed sum in (1.1) admits a decomposition into a simple combinatorial sum and another double-indexed sum induced by a degenerate tensor $\mathbf{D}=[d(i,j,k,\ell)]\in\mathbb{R}^{N\times N\times N\times N}$ (Zhao et al., 1997; Barbour and Chen, 2005), where degenerate means that all partial averages vanish; see section 2 for formal definition. Specifically, define
@@
In this paper, we develop new decoupling and randomization inequalities to simplify the task of bounding the moment-generating function (MGF) of $Q_{d}$, $\mathrm{E}\big[\exp(\lambda Q_{d})\big]$. Based on the bounds on the MGF of $Q_{d}$, we derive concentration inequalities for $Q_{d}$, and consequently also for the original statistic $Q_{w}$.
@@
where $\tau$ is an independent copy of $\pi$. The decoupling inequalities of interest compare $\mathrm{E}[\exp(\lambda Q_{d})]$ with (a variant of) $\mathrm{E}[\exp(\lambda\widetilde{Q}_{d})]$, which is typically much easier to analyze.
@@ 1.1 Motivation
Statistics of the form (1.1) are prevalent in nonparametric statistics and have also received increasing attention in causal inference. In nonparametric statistics, it has been noted that many quantities of statistical interest—including the correlation coefficients of Pearson, Kendall, Spearman, and more recently Chatterjee (Chatterjee, 2021), as well as discrepancy measures such as the Mann-Whitney-Wilcoxon statistic—can be expressed in this form and thus benefit from a unified treatment. In survey sampling and causal inference, such statistics have also become increasingly important within the framework of finite-population inference (Lei and Ding, 2021; Han, 2024), where they play a role analogous to that of U-statistics in the super-population framework.
In the probability and statistical theory literature, (1.1) may be viewed as natural higher-order extensions of Hoeffding’s statistics in the context of sampling without replacement (Hoeffding, 1951). The latter corresponds to the special—and in fact single-indexed—case $d(i,j,k,\ell)=a(i,k)$ for some matrix $\mathbf{A}=[a(i,k)]\in\mathbb{R}^{N\times N}$.
@@ 1.2 Generalized combinatorial quadratic forms
Our version of decoupling and randomization enables the following tail bound for the generalized combinatorial quadratic form,
@@
which coincides with the matrix correlation statistic of Barbour and Chen (2005).
@@
Here $\nu$ measures the largest possible absolute inner product obtainable by selecting half of the coordinates from a row of $\mathbf{A}$ and half of the coordinates from a row of $\mathbf{C}$, and then matching the selected coordinates via an arbitrary bijection. $\lfloor N/2\rfloor$ arises from the decoupling step in the proof, where the index set is split into complementary blocks of sizes $\lceil N/2\rceil$ and $\lfloor N/2\rfloor$. The maximization over $\sigma$ and $\widetilde{\sigma}$ parametrizes all possible ways of selecting and matching these coordinates.
@@ 1.3 Examples in nonparametric statistics
$\mathrm{P}\Big(Q_{\sf U}-\mathrm{E}[Q_{\sf U}]>t\Big)\leq\exp\Big(-\frac{K\,t^{2}}{\,Nmn+Nt\,}\Big)\;+\;\exp\Big(-\frac{K\,t^{2}}{\,\min\{m,n\}^{2}+\min\{m,n\}t\,}\Big).$
@@
$R$ has the same distribution as
@@
Daniels (1944)’s framework encompasses several classical statistics as special cases, including Pearson’s correlation, Spearman’s rank correlation $\rho$, and Kendall’s $\tau$, under suitable choices of score functions:
@@
Chatterjee (2021) has recently introduced a rank correlation as an appealing alternative to the correlation coefficients in Example 1.2; see, e.g., Dette et al. (2013), Han (2021), Chatterjee (2024), and Lin and Han (2022) for more discussions. Under (1.3), Chatterjee’s rank correlation has the same distribution as
@@
$\mathrm{P}\big(\xi_{N}-\mathrm{E}[\xi_{N}]\geq t\big)\leq\exp\Big(-\frac{KN^{2}t^{2}}{1+Nt}\Big)\;+\;\exp\Big(-\frac{KNt^{2}}{1+\,t}\Big).$
@@ 1.4 An example in causal inference
Regression adjustment is widely used to improve efficiency in randomized experiments. Lei and Ding (2021) studied this method in completely randomized experiments with a diverging number of covariates and derived a decomposition of the regression-adjusted estimator $\widehat{\tau}_{\mathrm{adj}}$ proposed by Lin (2013). In the decomposition, the bias terms appear in quadratic forms.
@@
be the set of units assigned to the treatment level $\omega$. The bias of $\widehat{\tau}_{\mathrm{adj}}$ can then be written as a function of
@@
Following Lei and Ding (2021), we assume $\mathbf{X}$ to be column-centered, which ensures that both $\mathbf{H}$ and $\mathbf{Q}(\omega)$ are doubly centered. Applying Corollary 2.1, we can show that there exists a universal constant $K>0$ such that, for any $t>0$,
@@ 2 Main results
$B:=\max_{\sigma\in\mathcal{S}_{N}}\Big\|[d(i,j,\sigma(i),\sigma(j))]_{i,j}\Big\|_{\sf op}~~{\rm and}~~V:=\frac{1}{N}\sum_{i,k\in[N]}\xi(i,k)^{2}+\frac{1}{N^{2}}\sum_{\begin{subarray}{c}i,j,k,\ell\in[N]\\ i\neq j,k\neq l\end{subarray}}d(i,j,k,\ell)^{2},$ (2.5)
@@
Recall $Q_{d}$ defined in (1.2). Then, for all $\lambda\in[0,1/5400B]$,
@@
Let $N\geq 20$ and $\mathbf{W}=[w(i,j,k,\ell)]\in\mathbb{R}^{N\times N\times N\times N}$ be a fixed tensor. Recall $Q_{w}$ defined in (1.1). Then, there exists a universal constant $K>0$ such that for any $t>0$,
@@
$\xi_{w}(i,k)=d_{w}(i,i,k,k)-\frac{1}{N}\sum_{k^{\prime}=1}^{N}d_{w}(i,i,k^{\prime},k^{\prime})-\frac{1}{N}\sum_{i^{\prime}=1}^{N}d_{w}(i^{\prime},i^{\prime},k,k)+\frac{1}{N^{2}}\sum_{i^{\prime},k^{\prime}\in[N]}d_{w}(i^{\prime},i^{\prime},k^{\prime},k^{\prime})$
${\rm and}~~V_{d}:=\frac{1}{N}\sum_{i,k\in[N]}\xi_{w}(i,k)^{2}+\frac{1}{N^{2}}\sum_{\begin{subarray}{c}i,j,k,\ell\in[N]\\ i\neq j,k\neq l\end{subarray}}d_{w}(i,j,k,\ell)^{2},$
with $a_{w}(\cdot,\cdot)$ and $d_{w}(\cdot,\cdot,\cdot,\cdot)$ introduced in (2) and (2), respectively.
@@ 3 Proofs of the main results
In the following, recall that $G_{1},G_{2},\ldots,G_{1}^{\prime},G_{2}^{\prime},\ldots,$ are independent standard Gaussian random variables, further independent of the system. Recall that $\bigl[\cdot\bigr]_{i,j}$ denotes the matrix whose $(i,j)$-th entry is given by the enclosed expression, where $i$ and $j$ range over the prescribed index sets (not necessarily $[N]$). For a set $I\subseteq[N]$, we write $|I|$ for its cardinality and $I^{c}:=[N]\setminus I$ for its complement. Additionally, for a permutation (or, more generally, a mapping) $\pi$, we denote $\pi(I):=\{\pi(i)\mid i\in I\}$. For a matrix $A=[a(i,j)]\in\mathbb{R}^{N\times N}$, define
$f(\pi):=\sum_{i=1}^{N}a(i,\pi(i)).$ (3.1)
When the dependence on $\pi$ needs to be emphasized, we write $f(\pi)$; otherwise, we write $f$ for simplicity.
@@ 3.1 Auxiliary lemmas
Let $\mathbf{A}=[a(i,j)]\in\mathbb{R}^{N\times N}$ be a (possibly random) matrix independent of $\pi$ and satisfying $\sum_{i,j\in[N]}a(i,j)=0$. We then have, for any $\lambda\in\mathbb{R}$,
@@
Let $\mathbf{A}=[a(i,j)]\in\mathbb{R}^{N\times N}$ be a fixed matrix. We then have, for any $|\lambda|<\sqrt{2}/B_{A}$,
@@
Let $\mathbf{A}=[a(i,j)]\in\mathbb{R}^{N\times N}$ be a fixed matrix whose entries are in $[0,1]$ and $f$ be defined as in (3.1). Then for any $\lambda\geq 0$,
@@
Let $\mathbf{A}=[a(i,j)]\in\mathbb{R}^{N\times N}$ be a fixed matrix whose entries are in [1,1][-1,1] and $f$ be defined as in (3.1). Then for any $|\lambda|\leq 1/3$,
@@
Fix $\mathbf{A}$ as in Lemma 3.6 and let $f$ be defined as in (3.1). Then for any $\lambda\geq 0$,
@@
$V_{2}=\max_{I,J}\frac{1}{N_{1}(N-N_{1})}\sum_{\begin{subarray}{c}i\in I,k\in J\\ j\in I^{c},\ell\in J^{c}\end{subarray}}\widetilde{d}_{I,J}(i,j,k,\ell)^{2}\leq\max_{I,J}\frac{1}{N_{1}(N-N_{1})}\sum_{\begin{subarray}{c}i\in I,k\in J\\ j\in I^{c},\ell\in J^{c}\end{subarray}}d_{I,J}(i,j,k,\ell)^{2}\leq 4.1V$
@@
$\xi_{w}(i,k)$ $=d_{w}(i,i,k,k)-\frac{1}{N}\sum_{k^{\prime}=1}^{N}d_{w}(i,i,k^{\prime},k^{\prime})-\frac{1}{N}\sum_{i^{\prime}=1}^{N}d_{w}(i^{\prime},i^{\prime},k,k)+\frac{1}{N^{2}}\sum_{i^{\prime},k^{\prime}\in[N]}d_{w}(i^{\prime},i^{\prime},k^{\prime},k^{\prime})$
$=\Big(c_{ii}-\frac{1}{N}\sum_{i^{\prime}=1}^{N}c_{i^{\prime}i^{\prime}}\Big)\Big(\widetilde{a}_{kk}-\frac{1}{N}\sum_{k^{\prime}=1}^{N}\widetilde{a}_{k^{\prime}k^{\prime}}\Big).$
@@ 4 Proofs of auxiliary lemmas
By Chatterjee’s third lemma Han (2024, Lemma 3.11), taking $\psi=\frac{3}{2}\lambda^{2}$, we have
@@
where $r(\cdot)$ is defined as in Han (2024, Lemma 3.11); for completeness, we recall that for any $\psi>0$,
@@
$f^{\downarrow}:=\sum_{i=1}^{N}a(i,\pi(i))\cdot\mathds{1}(|a(i,\pi(i))|\leq\rho)~~{\rm and}~~f^{\uparrow}:=\sum_{i=1}^{N}|a(i,\pi(i))|\cdot\mathds{1}(|a(i,\pi(i))|>\rho),$